# CDsampling: An R Package for Constrained D-Optimal Sampling in Paid Research Studies

### v2

@@ 3.2   Applications in paid research study: $\tt trauma\_data$ example
> nsample=600 #sample size> constraint = c(392, 410)  #mild:severe> lower.bound <- function(i, w0){+   n = 600+   constraint = c(392,410)+   if(i <= 4){+     a.lower <- (sum(w0[5:8])-(constraint[2]/n)*(1-w0[i]))/(sum(w0[5:8]))}+   else{+     a.lower <- (sum(w0[1:4])-(constraint[1]/n)*(1-w0[i]))/(sum(w0[1:4]))}+    a.lower}> upper.bound <- function(i, w0){+  n = 600+  constraint = c(392,410)+  if(i <= 4){+    b.upper <- ((constraint[1]/n)*(1-w0[i]) - (sum(w0[1:4])-w0[i]))/(1-sum(w0[1:4]))}+  else{+    b.upper <- ((constraint[2]/n)*(1-w0[i]) - (sum(w0[5:8])-w0[i]))/(1-sum(w0[5:8]))}+    b.upper}> g.con = matrix(0,nrow=length(constraint)+1+m, ncol=m)> g.con[2:3,] = matrix(data=c(1,1,1,1,0,0,0,0,0,0,0,0,1,1,1,1), ncol = m, byrow=TRUE)> g.con[1,] = rep(1, m)> g.con[4:(length(constraint)+1+m), ] = diag(1, nrow=m)> g.dir = c("==", "<=","<=", rep(">=",m))> g.rhs = c(1, ifelse((constraint/nsample<1),constraint/nsample,1), rep(0, m))
@@
We then run the constrained lift-one algorithm to find the constrained D-optimal approximate allocation using liftone_constrained_MLM() function and then convert the approximate allocation to an exact allocation with approxtoexact_constrained_func function.
@@
> set.seed(123)> approx_design = liftone_constrained_MLM(m=m, p=p, Xi=Xi, J=J, beta=beta,+ lower.bound=lower.bound, upper.bound=upper.bound, g.con=g.con, g.dir=g.dir,+ g.rhs=g.rhs, w00=NULL, link=’cumulative’,  Fi.func=Fi_func_MLM, reltol=1e-5,+ maxit=500, delta=1e-6, epsilon=1e-8, random=TRUE, nram=3)> exact_design = approxtoexact_constrained_func(n=600, w=approx_design\$w, m=8,+ beta=beta, link=’cumulative’, X=Xi, Fdet_func=Fdet_func_MLM,+ iset_func=iset_func_trauma)> print(exact_design)Optimal Sampling Results:================================================================================allocation :155.0, 0.0, 0.0, 100.0, 168.0, 0.0, 0.0, 177.0--------------------------------------------------------------------------------allocation.real :0.2593, 0.0, 0.0, 0.1667, 0.2796, 0.0, 0.0, 0.2944--------------------------------------------------------------------------------det.maximum :1.63163827059162e+23--------------------------------------------------------------------------------
@@
As the trauma_data example doesn’t have bounded constraints, to find the constrained uniform sampling allocation, we use approxtoexact_constrained_func() with one subject in each stratum or category as the starting point, that is, the approximate allocation $\mathbf{w}=(1/600,1/600,1/600,1/600,$ 1/600,1/600,1/600,1/600)1/600,1/600,1/600,1/600)1 / 600 , 1 / 600 , 1 / 600 , 1 / 600 ). The corresponding $I$ set function is provided in the CDsampling package, and it can be easily defined according to other constraints. Note that the determinant provided by approxtoexact_constrained_func() for different designs are not comparable, as the criteria Fdet_func differ.
@@
> unif_design = approxtoexact_constrained_func(n=600, w=rep(1/600,8), m=8,+ beta=NULL, link=NULL, X=NULL, Fdet_func=Fdet_func_unif, iset_func=iset_func_trauma)> print(unif_design)Optimal Sampling Results:================================================================================allocation :75.0, 75.0, 75.0, 75.0, 75.0, 75.0, 75.0, 75.0--------------------------------------------------------------------------------allocation.real :0.0017, 0.0017, 0.0017, 0.0017, 0.0017, 0.0017, 0.0017, 0.0017--------------------------------------------------------------------------------det.maximum :1001129150390625--------------------------------------------------------------------------------
@@ 4   Summary
The current version of CDsampling implements D-optimal allocations within both paid research sampling and general study frameworks with or without constraints. Its primary objective is to optimize sampling allocations for better model estimation accuracy in the studies. The package includes F_func_GLM() and F_func_MLM() for the computation of the Fisher information matrix of GLMs and MLMs, respectively. It is noteworthy that standard linear regression models are special GLM with an identity link function and Gaussian-distributed responses, which is also supported by our package. Illustrative examples are provided in Supplementary Section S2.
To find standard or unconstrained D-optimal allocations, our package implements the lift-one algorithm through functions liftone_GLM() and liftone_MLM(). Paid research studies often impose sampling constraints. To address this, the constrained lift-one algorithm can be applied using functions liftone_constrained_GLM() and liftone_constrained_MLM(). An example illustrating the difference between the lift-one algorithm and the constrained lift-one algorithm is provided in Supplementary Section S1.
@@
S1 Comparison of lift-one algorithm and constrained lift-one algorithm: An example of comparison between lift-one algorithm and constrained lift-one algorithm;
S2 Examples of Fisher information matrix: Two examples of finding Fisher information matrix with CDsampling;
@@
S1   Comparison of lift-one algorithm and constrained lift-one algorithm
@@
> w00 = c(1/6, 1/6, 2/3) #an arbitrary starting allocation> liftone_GLM(X=X, W=W_matrix, reltol=1e-10, maxit=100, random=FALSE, nram=3, w00=w00)Optimal Sampling Results:================================================================================w :0.3333, 0.3333, 0.3333--------------------------------------------------------------------------------w0 :0.1667, 0.1667, 0.6667--------------------------------------------------------------------------------Maximum :0.0077--------------------------------------------------------------------------------itmax :8.0--------------------------------------------------------------------------------convergence :TRUE--------------------------------------------------------------------------------
@@
We consider the same logistic regression model (S2.1) as in Examples S2.1 and S1.1. In this example, we aim for the constrained D-optimal allocation in $S=\{(w_{1},w_{2},w_{3})^{\top}\in S_{0}\mid w_{1}\leq\frac{1}{6},w_{3}\geq\frac{8}{15},4w_{1}\geq w_{3}\}$. The additional constraints on $S_{0}$ can be defined using the following R codes:
@@
We also need to find the upper boundary and lower boundary functions to find $[r_{i1},r_{i2}]$ in step $3^{\circ}$ of the constrained lift-one algorithm to meet the conditions in $S$. The $r_{i1}$ and $r_{i2}$ can be obtained as follows:
@@
Then we are ready to find the constrained D-optimal allocation in $S$ using the constrained lift-one function liftone_constrained_GLM().
> set.seed(123)> liftone_constrained_GLM(X=X, W=W_matrix, g.con=g.con, g.dir=g.dir, g.rhs=g.rhs,+ lower.bound=lower.bound, upper.bound=upper.bound, reltol=1e-10, maxit=100,+ random=FALSE, nram=3, w00=w00, epsilon=1e-8)Optimal Sampling Results:================================================================================w :0.1667, 0.3, 0.5333--------------------------------------------------------------------------------w0 :0.1667, 0.1667, 0.6667--------------------------------------------------------------------------------maximum :0.0055--------------------------------------------------------------------------------convergence :TRUE--------------------------------------------------------------------------------itmax :1.0--------------------------------------------------------------------------------deriv.ans :0.0199, 0.0026, -0.0133--------------------------------------------------------------------------------gmax :0.0--------------------------------------------------------------------------------reason :"gmax <= 0"--------------------------------------------------------------------------------
@@
S2   Examples of the Fisher information matrix
In Example S2.1 below, we provide an example for calculating a GLM Fisher information matrix through the CDsampling package.
Example S2.1.
Consider a research study under a logistic regression model
We also provide in Example S2.2 with the computation of the MLM Fisher information matrix through the CDsampling package.
Example S2.2.
Consider a research study under a cumulative logit npo model with $J=5$ response levels and covariates $(x_{i1},x_{i2})\in\{(1,0),(2,0),(3,0),(4,0),(1,1),(2,1),(3,1),$ (4,1)}(4,1)\}( 4 , 1 ) }. The model can be written as the following four equations:
To calculate the Fisher information matrix $\mathbf{F}$ for the MLM in this example, we use the function F_func_MLM() in the CDsampling package. In the R code below, $J$ is the number of levels of the multinomial response, $p$ is the number of parameters, $m$ is the number of experimental settings or covariate combinations, and $X_{i}$ is the model matrix at the $i$th experimental setting. The function F_func_MLM() returns the Fisher information matrix ${\mathbf{F}}({\mathbf{w}})$ given ${\mathbf{w}}$.
$\Box$

### v3

@@ 3.2   Applications in paid research study: $\tt trauma\_data$ example
> nsample=600 #sample size> constraint = c(392, 410)  #mild:severe> #lower bound function in step 3 of constrained liftone> lower.bound <- function(i, w0){+   n = 600+   constraint = c(392,410)+   if(i <= 4){+     a.lower <- (sum(w0[5:8])-(constraint[2]/n)*(1-w0[i]))/(sum(w0[5:8]))}+   else{+     a.lower <- (sum(w0[1:4])-(constraint[1]/n)*(1-w0[i]))/(sum(w0[1:4]))}+    a.lower}> #upper bound function in step 3 of constrained liftone> upper.bound <- function(i, w0){+  n = 600+  constraint = c(392,410)+  if(i <= 4){+    b.upper <- ((constraint[1]/n)*(1-w0[i]) - (sum(w0[1:4])-w0[i]))/(1-sum(w0[1:4]))}+  else{+    b.upper <- ((constraint[2]/n)*(1-w0[i]) - (sum(w0[5:8])-w0[i]))/(1-sum(w0[5:8]))}+    b.upper}> #define constraints> g.con = matrix(0,nrow=length(constraint)+1+m, ncol=m)> g.con[2:3,] = matrix(data=c(1,1,1,1,0,0,0,0,0,0,0,0,1,1,1,1), ncol = m, byrow=TRUE)> g.con[1,] = rep(1, m)> g.con[4:(length(constraint)+1+m), ] = diag(1, nrow=m)> g.dir = c("==", "<=","<=", rep(">=",m))> g.rhs = c(1, ifelse((constraint/nsample<1),constraint/nsample,1), rep(0, m))
Then, we may define an optional label of the sampling subgroups that corresponds to each of the $m=8$ subgroups using the following code:
> label=label = c("Placebo-Mild", "Low-Mild", "Medium-Mild", "High-Mild", "Placebo-Severe",+ "Low-Severe", "Medium-Severe", "High-Severe")
@@
We then run the constrained lift-one algorithm to find the constrained D-optimal approximate allocation using liftone_constrained_MLM() function and convert the approximate allocation to an exact allocation with approxtoexact_constrained_func function.
@@
> set.seed(123)> approx_design = liftone_constrained_MLM(m=m, p=p, Xi=Xi, J=J, beta=beta,+ lower.bound=lower.bound, upper.bound=upper.bound, g.con=g.con, g.dir=g.dir,+ g.rhs=g.rhs, w00=NULL, link=’cumulative’,  Fi.func=Fi_func_MLM, reltol=1e-5,+ maxit=500, delta=1e-6, epsilon=1e-8, random=TRUE, nram=3, label=label)> exact_design = approxtoexact_constrained_func(n=600, w=approx_design\$w, m=8,+ beta=beta, link=’cumulative’, X=Xi, Fdet_func=Fdet_func_MLM,+ iset_func=iset_func_trauma, label=label)> print(exact_design)Optimal Sampling Results:================================================================================Optimal exact allocation:             Placebo-Mild Low-Mild Medium-Mild High-Mild Placebo-Severeallocation      155.0        0.0      0.0         100.0     168.0allocation.real 0.2593       0.0      0.0         0.1667    0.2796              Low-Severe Medium-Severe High-Severeallocation      0.0        0.0           177.0allocation.real 0.0        0.0           0.2944--------------------------------------------------------------------------------det.maximum :1.63163827059162e+23--------------------------------------------------------------------------------
The allocation output provides the exact allocation of the sampling across different treatment-severity subgroups, representing the implementable sample sizes for each subgroup. The result is derived by converting the allocation.real, which is the D-optimal approximate allocation outcome from liftone_constrained_MLM().
@@
As the trauma_data example doesn’t have bounded constraints, to find the constrained uniform sampling allocation, we use approxtoexact_constrained_func() with one subject in each stratum or subgroup as the input, that is, the approximate allocation $\mathbf{w}=(1/600,1/600,1/600,1/600,$ 1/600,1/600,1/600,1/600)1/600,1/600,1/600,1/600)1 / 600 , 1 / 600 , 1 / 600 , 1 / 600 ). The corresponding $I$ set function is provided in the CDsampling package, and it can be easily defined according to other constraints, see Section S4 of Supplementary Material. Note that the determinant provided by approxtoexact_constrained_func() for different designs are not comparable, as the criteria Fdet_func differ.
@@
> unif_design = approxtoexact_constrained_func(n=600, w=rep(1/600,8), m=8,+ beta=NULL, link=NULL, X=NULL, Fdet_func=Fdet_func_unif, iset_func=iset_func_trauma)> print(unif_design)Optimal Sampling Results:================================================================================Optimal exact allocation:                Placebo-Mild Low-Mild Medium-Mild High-Mild Placebo-Severe Low-Severeallocation      75.0         75.0     75.0        75.0      75.0           75.0allocation.real 0.0017       0.0017   0.0017      0.0017    0.0017         0.0017                Medium-Severe High-Severeallocation      75.0          75.0allocation.real 0.0017        0.0017--------------------------------------------------------------------------------det.maximum :1001129150390625--------------------------------------------------------------------------------
@@ 4   Summary
The current version of CDsampling implements D-optimal allocations within both paid research sampling and general study frameworks with or without constraints. Its primary objective is to optimize sampling allocations for better model estimation accuracy in the studies. The package includes F_func_GLM() and F_func_MLM() for the computation of the Fisher information matrix of GLMs and MLMs, respectively. It is noteworthy that standard linear regression models are special GLM with an identity link function and Gaussian-distributed responses, which is also supported by our package. Theoretical results are summarized in Section 2.2 and Section 2.3 while illustrative examples are provided in Supplementary Section S1.
To find standard or unconstrained D-optimal allocations, our package implements the lift-one algorithm through functions liftone_GLM() and liftone_MLM(). Paid research studies often impose sampling constraints. To address this, the constrained lift-one algorithm can be applied using functions liftone_constrained_GLM() and liftone_constrained_MLM(). An example illustrating the difference between the lift-one algorithm and the constrained lift-one algorithm is provided in Supplementary Section S2 while Section 3 presents two application examples from paid research studies.
@@
S1 Examples of Fisher information matrix: Two examples of finding Fisher information matrix with CDsampling;
S2 Comparison of lift-one algorithm and constrained lift-one algorithm: An example of comparison between lift-one algorithm and constrained lift-one algorithm;
@@
S4 Template functions of constraint index set $I$: Detailed R codes for the example-specific functions of constraint index set $I$ in CDsampling;
S1   Examples of the Fisher information matrix
In Example S1.1 below, we provide an example of calculating the Fisher information matrix for GLM through the CDsampling package.
Example S1.1.
Consider a research study under a logistic regression model
We also provide in Example S1.2 with the computation of the Fisher information matrix for MLM through the CDsampling package.
Example S1.2.
We revisit the trauma_data example previously analyzed in Section 3.2. Consider a research study under a cumulative logit npo model with $J=5$ response levels and covariates $(x_{i1},x_{i2})\in\{(1,0),(2,0),(3,0),(4,0),(1,1),(2,1),(3,1),$ (4,1)}(4,1)\}( 4 , 1 ) }. The model can be written as the following four equations:
To calculate the Fisher information matrix $\mathbf{F}$ for the MLM in this example, we use the function F_func_MLM() in the CDsampling package. In the R code below, $J$ is the number of levels of the response variable, $p$ is the number of parameters in the model, $m$ is the number of design points or subgroups, and $X_{i}$ is the model matrix of the $i$th design point. The function F_func_MLM() returns the Fisher information matrix ${\mathbf{F}}({\mathbf{w}})$ given ${\mathbf{w}}$.
$\Box$
S2   Comparison of lift-one algorithm and constrained lift-one algorithm
@@
> w00 = c(1/6, 1/6, 2/3) #an arbitrary starting allocation> liftone_GLM(X=X, W=W_matrix, reltol=1e-10, maxit=100, random=FALSE, nram=3, w00=w00)Optimal Sampling Results:================================================================================Optimal approximate allocation:   1      2      3w  0.3333 0.3333 0.3333w0 0.1667 0.1667 0.6667--------------------------------------------------------------------------------Maximum :0.0077--------------------------------------------------------------------------------itmax :7.0--------------------------------------------------------------------------------convergence :TRUE--------------------------------------------------------------------------------
@@
We consider the same logistic regression model (S1.1) as in Examples S1.1 and S2.1. In this example, we aim for the constrained D-optimal allocation in $S=\{(w_{1},w_{2},w_{3})^{\top}\in S_{0}\mid w_{1}\leq\frac{1}{6},w_{3}\geq\frac{8}{15},4w_{1}\geq w_{3}\}$. The additional constraints on $S_{0}$ can be defined using the following R codes:
@@
We also need to find the upper boundary and lower boundary functions for $[r_{i1},r_{i2}]$ in step $3^{\circ}$ of the constrained lift-one algorithm to meet the conditions in $S$. The $r_{i1}$ and $r_{i2}$ can be obtained as follows:
@@
Now we are ready to find the constrained D-optimal allocation in $S$ using the constrained lift-one function liftone_constrained_GLM().
> set.seed(123)> liftone_constrained_GLM(X=X, W=W_matrix, g.con=g.con, g.dir=g.dir, g.rhs=g.rhs,+ lower.bound=lower.bound, upper.bound=upper.bound, reltol=1e-10, maxit=100,+ random=FALSE, nram=3, w00=w00, epsilon=1e-8)Optimal Sampling Results:================================================================================Optimal approximate allocation:   1      2      3w  0.1667 0.3    0.5333w0 0.1667 0.1667 0.6667--------------------------------------------------------------------------------maximum :0.0055--------------------------------------------------------------------------------convergence :TRUE--------------------------------------------------------------------------------itmax :1.0--------------------------------------------------------------------------------deriv.ans :0.0199, 0.0026, -0.0133--------------------------------------------------------------------------------gmax :0.0--------------------------------------------------------------------------------reason :"gmax <= 0"--------------------------------------------------------------------------------
@@
S4   Template functions of constraint index set $I$
> iset_func_trauma <- function(allocation){+   iset = rep(TRUE,8)+   if(sum(allocation[1:4])>=392){iset[1:4]=FALSE}+   if(sum(allocation[5:8])>=410){iset[5:8]=FALSE}+   return(iset)+ }turn(iset)}
> iset_func_trial <- function(allocation){+   Ni = c(50, 40, 10, 200, 150, 50)+   return(allocation < Ni)+ }